At a toy factory, each worker assembles one toy at a time and can assemble a total of 48 toys in 2 hours. At that rate, how many toys can 4 workers assemble in 3 hours?
Answer by ankor@dixie-net.com(22740)

each worker assembles one toy at a time and can assemble a total of 48 toys in 2 hours.
then
48/2 = 24 toys per man-hour
:
At that rate, how many toys can 4 workers assemble in 3 hours?
4 working for 3 hrs: 4*3 = 12 man-hours
then
12 * 24 = 288 toys